OpenStudy (anonymous):

What is the graph of the function rule? y = |4x| + 3

OpenStudy (agent0smith):

You could make a table of values if you're clueless on how to graph it. Or, graph it using transformations of |x|.

OpenStudy (anonymous):

it's a peace sign. LOL v(^^)v

OpenStudy (anonymous):

Well I know I have to subtract 3, so I can get 4x by itself. But what else do I subtract by 3 if we don't know what y is?

OpenStudy (anonymous):

@math&ing001 @mathmale

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

I think the function rules is when the variables are related, you can make a table

OpenStudy (the_fizicx99):

I think you should make a table an replace it with r x values. Ex: -2 and plug that in and solve

OpenStudy (the_fizicx99):

x value's *

OpenStudy (the_fizicx99):

f(x) is a fancy way of saying y. For example here:

OpenStudy (the_fizicx99):

Can't give you the answer I already got a warning today for no reason..

OpenStudy (agent0smith):

"Well I know I have to subtract 3, so I can get 4x by itself. But what else do I subtract by 3 if we don't know what y is?" Do not do this, you aren't solving for x. Make a table of x and y values.

OpenStudy (anonymous):

I'm thinking it's this one: http://prntscr.com/2cnbv5

OpenStudy (the_fizicx99):

Make a table, plot the point and draw the line, simple :)

OpenStudy (agent0smith):

Are you sure it's that one? Describe why you think it's that one.

OpenStudy (anonymous):

Well because it adds three... And it seems to start as a positive. Ugh I'm so bad at this.

OpenStudy (anonymous):

If it's 4x + 3 It can be like 4(1) + 3 = 7 4(2) + 3 = 11 4(3) + 3 = 15 Is it like this..?

OpenStudy (the_fizicx99):

I'd include a negative, -2 -1 0 1 2

OpenStudy (anonymous):

4(-2) + 3 = -5 4(-1) + 3 = -1 4(0) + 3 = 3 4(1) + 3 = 7 4(2) + 3 = 11 Like this?

OpenStudy (agent0smith):

You ignored the absolute value sign.

OpenStudy (anonymous):

Oh true. So that would make the -2 and the -1, positive 2 and 1, right?

OpenStudy (the_fizicx99):

Yesh

OpenStudy (anonymous):

Because it would be |4(-2)| and that would equal |-8| and that would actually be 8? And then add the 3, of course.

OpenStudy (agent0smith):

Yes.

OpenStudy (the_fizicx99):

Did that answer you're question? :)

OpenStudy (anonymous):

Yes but I don't understand how I'd graph that.

OpenStudy (the_fizicx99):

Plot the points? and draw the line

OpenStudy (agent0smith):

That's the point of a table of values. You write out the table of x and y, so you can plot the points.

OpenStudy (anonymous):

Just to make sure... |4(-2)| + 3 = 11.... That would be (-2, 11), right..? If x, in the problem, was -2.

OpenStudy (agent0smith):

Yes, that's why you should write a table to avoid any confusion.

OpenStudy (the_fizicx99):

Yup

OpenStudy (anonymous):

Okay, yeah, now I understand it.

OpenStudy (anonymous):

Thank you guys so much, sorry it took me a while. Been a long week ._.

OpenStudy (the_fizicx99):

Nw

OpenStudy (anonymous):

This is what I got: http://prntscr.com/2cnbv5

OpenStudy (agent0smith):

Yes.

OpenStudy (the_fizicx99):

yes